OpenStudy (anonymous):

Write an equation in slope-intercept form for the line with a slope m= 2/3 and a y-intercept b= -10.

OpenStudy (midhun.madhu1987):

Slope-Intercept form of a line is given by y = mx + c where m = slope and c = y-intercept

OpenStudy (midhun.madhu1987):

Plug-in the values...

OpenStudy (anonymous):

can you please show me

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

m = 2/3 and c = -10

OpenStudy (anonymous):

so the equation would be y=2/3 -10 ?

OpenStudy (midhun.madhu1987):